Encyclopedia Astronautica
2008.06.05 - ISS On-Orbit Status 06/05/08


Flight Day 6 (FD6) of STS-124/1J.

ISS crew work cycle remains unchanged: wake 6:32am EDT; sleep 10:02pm.

Crew activities aboard the ISS today centered on three major areas: (1) The second 1J spacewalk, (2) activation of the JPM (Japanese Pressurized Module) Kibo, and (3) preparations for tomorrow's relocation of the JLP (Japanese Logistics Pressurized Module).

Mission 1J's EVA-2 was completed successfully by Mike Fossum & Ron Garan in 7h 11min, accomplishing all its objectives. (During the spacewalk, Fossum (EV1) & Garan (EV2) -

Installed Fwd/Aft JTVEs (JEM Television Equipment, i.e., 2 cameras) for monitoring JEM RMS ops (~12:19pm),
Removed 7 thermal covers from each of the JEM RMS' six joints & one end effector (~1:02pm),
Prepared the JPM zenith ACBM (Active Common Berthing Mechanism) for JLP relocation on FD7 (removed ACBM cover; released MMOD/Micrometeoroid Orbital Debris shield launch restraints),
Cleared up late-discovered MLI (Multi-Layer Insulation) cover anomaly on the JPM zenith CBM (could have interfered with the 4 capture latches during JLP relocation/berthing tomorrow) (~1:55pm),
Installed JPM trunnion and keel pin thermal covers (~2:15pm),
Prepared the old S1 Nitrogen Tank Assembly on S1 truss and new spare on ESP-3 (External Stowage Platform 3) for R&R during EVA 3 (~3:45pm), and
Retrieved an ETVCG (External TV Camera Group) camera from CP9 (Camera Port 9) on the left truss.
Official start time of the spacewalk was 11:04am EDT, about 28 minutes ahead of timeline, and it ended at 6:15pm. Total EVA duration (PET = Phase Elapsed Time) was 7h 11min. It was the 110th spacewalk for ISS assembly & maintenance and the 82nd from the station (60 from Quest, 22 from Pirs, plus 28 from Shuttle) totaling 506h 46min, the second for Expedition 17 and the 11th so far this year. After today's EVA, a total of 141 spacewalkers (109 NASA astronauts, 21 Russians, and 11 astronauts representing Japan-1, Canada-4, France-1, Germany-2 and Sweden-3) have logged a total of 695h 08min outside the station on building, outfitting & servicing. It was the 132nd spacewalk involving U.S. astronauts.)

After wakeup at ~6:32am, ending the 8.5-hr sleep period before the spacewalk, the A/L CL (Airlock Crewlock) hatch was cracked (~7:12am) for a hygiene break/with mask prebreathe for Fossum & Garan, after spending the night on 10.2 psi campout. Around 8:22am, the hatch was closed again by IVs (Intravehicular Crewmembers) for EVA preparations in 10.2 psi, followed by EMU (Extravehicular Mobility Unit) purge & prebreathe. Sergey Volkov, Mark Kelly & Greg Chamitoff assisted the spacewalkers during EMU purge, EMU prebreathe and CL depressurization. EV1/EV2 egress followed after the EMUs were switched to batteries at 11:04am.

Before breakfast, Chamitoff collected his second 'wet' saliva sample for the biomed experiment INTEGRATED IMMUNE (Validating Procedures for Monitoring Crew member Immune Function). (IMMUNE protocol requires the collection to occur first thing post-sleep, before eating, drinking and brushing teeth, and all samples are stored at ambient temperature. Along with NUTRITION (Nutritional Status Assessment), INTEGRATED IMMUNE samples & analyzes participant's blood, urine, and saliva before, during and after flight for changes related to functions like bone metabolism, oxidative damage and immune function to develop and validate an immune monitoring strategy consistent with operational flight requirements and constraints. The strategy uses both long and short duration crewmembers as study subjects. The saliva is collected in two forms, dry and liquid. The dry samples are collected at intervals during the collection day using a specialized book that contains filter paper. The liquid saliva collections require that the crewmember soak a piece of cotton inside their mouth and place it in a salivette bag; there are four of the liquid collections during docked operations.)

Prior to EVA start, CDR Volkov activated two Kodak DCS760 digital cameras with 28mm lens for the EVA and verified deactivation of all onboard ham radio equipment (Kenwood in SM/Service Module, Ericsson in FGB) to prevent RF interference with the EMUs, as well as proper closure of the protective Lab window shutters.

As the spacewalk went on, monitored by Ken Ham & Mark Kelly, crewmembers Hoshide, Nyberg, Reisman, Volkov, Kononenko & Chamitoff worked in the JPM at various times on carefully choreographed outfitting of Kibo. Major transfers from JLP and commensurate installation activities in JPM included -

EPS1 (Electrical Power Systems 1) rack & mating of its umbilicals,
DMS1 (Data Management Systems 1) rack & mating of its umbilicals,
WS (Workstation) rack and checking out its C&W (Caution & Warning) panel,
ICS (Interorbit Communication System) rack,
SAIBO ('biologic cell') payload rack after installing two monitor display covers (CB/Clean Bench; CBEF/Cell Biology Experiment Facility),
RYUTAI ('fluid') payload rack,
JRSR (JEM Resupply Stowage) rack,
Mating umbilicals for the JEMRMS (Japanese Experiment Module Robotic Manipulator System) console rack transferred last night from the JLP, checking out its C&W (Caution & Warning) panel, and configuring its console (deploying two cue cards),
Removing launch restraints of the pre-installed EPS2 & DMS racks, and
Relocating two dummy hard panels from JPM to locations in JLP.
For the JPM outfitting, Hoshide & Nyberg also -

Deployed the remaining Photo/TV drag-thru cable from Node-2,
Reconfigured both the 'B' string & 'A' string HCTLs (Heater Controllers) from LTA (Launch-to-Activation) to On-Orbit, reconnecting appropriate power cables,
Installed the JPM Channel 'A' power jumper (W3009) in the Vestibule (after transfer & connection of EPS1 & DMS1), allowing TKSC (Tsukuba Space Center)/Japan to assume Kibo control,
Moved the JPM PCS (Portable Computer System) laptop from its temporary location to the JEMRMS rack and mounted it on the rack's right side,
Set up two SSC (Station Support Computer) laptops in the JPM, one hard-wired, the other off a drag-thru cable (to enable Japanese language, IP Phone, etc.), and
Configured the DOUG (Dynamic Onboard Ubiquitous Graphics) application on the PCS for the upcoming activation of the RMS in front of Kibo, to receive & display SRMS/SSRMS/JEMRMS joint angle telemetry from the laptop. (DOUG is a software application that provides a graphical birdseye-view image of the external station configuration and the JEMRMS, showing its real-time location and configuration on a laptop during its operation.)
In preparation for tomorrow's relocation of the JLP module from Node-2 zenith to JPM with the SSRMS (Space Station Remote Manipulator System) --

Nyberg deactivated the JLP MKAM (Minimum Keep Alive Monitor) fan which had been turned on temporarily by Reisman earlier this morning as usual to allow ingress into JLP,
Aki Hoshide unlatched the JPM's overhead hatch,
Garrett Reisman installed & checked out the CBCS (Centerline Berthing Camera System) at the JPM hatch, and
Karen & Oleg installed the CPA (Controller Panel Assembly) at the Node-2 zenith hatch so that Flight Controllers at MCC-Houston can make preparations for the JLP relocation during crew sleep,
Sergey walked off the SSRMS, i.e., ungrappled the JPM PDGF (Power & Data Grapple Fixture), which had provided vital keep-alive power to the JPM, and maneuvered the arm to grapple, at 6:08pm, the MBS (Mobile Base System) PDGF-3 for a ground-commanded base change later tonight (the MT/Mobile Transporter with the MBS was moved yesterday from WS4/Worksite 4 to WS6).
In the US Lab, MS-1 Nyberg installed the AmiA (Antimicrobial Applicator), brought up on STS-124, in the ITCS (Internal Thermal Control System) and removed its insulation blanket to allow temperature equalization. Later tonight, AmiA will be removed again. (AmiA is introducing OPA (Ortho-Phthalaldehyde), an antimicrobial agent, into the Lab ITCS coolant.)

Sergey Volkov terminated battery charging for the 'Kelvin-Video' and TTM-2 instruments for another operational run of the Russian KPT-2 science payload BAR-RM, starting tomorrow (6/6) TTM-2 using the RSE-1 laptop, with downlinking via BSR-TM channel. (Objective of the payload is to experiment with ISS leak detection based on environmental data anomalies (temperature, humidity, and ultrasound emissions) at leak locations. The payload uses a remote infrared thermometer (Kelvin-Video), a thermohygrometer (Iva-6A), a heat-loss anemometer/thermometer (TTM-2), an ultrasound analyzer (AU-01), and a leak detector (UT2-03) to determine physical background signs of loss of ISS pressure integrity which could be indicative of leaks in the working compartments of the station. Measurements are taken in specific zones (13 in SM PkhO and 4 in DC1), both with lights & fans turned on and off. )

The FE-1 conducted the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS (Russian Segment) hatchways, including the passageways PrK (SM Transfer Compartment) -ATV, PrK -RO (SM Working Compartment), PkhO (SM Transfer Tunnel) -RO, PkhO -DC1, PkhO -FGB PGO, FGB PGO -FGB GA, FGB GA -Node-1. (This checkup is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).)

In preparation for a major RS outfitting activity scheduled tomorrow, involving the nominal restoration of mounting structures for the heat exchanger-condenser fans VTK1 & VTK2 of the two Russian SKV air conditioners, the CDR unstowed and readied necessary tooling and reviewed procedures.

After the spacewalkers' return on board at 6:15pm, post-EVA activities by Fossum, Garan, Kelly and Chamitoff in the A/L consist of --

Recharging the EMU/spacesuits with water from PWR (Payload Water Reservoir),
Reconnecting the LTAs (Lower Torso Assemblies) to the EMUs,
Capping the UIA (Umbilical Interface Assembly (no METOX regeneration required),
Setting up the METOX (Metal Oxide) CO2 absorbers for regeneration in the A/L bake-out oven,
Downlinked EVA imagery from the Kodak DCS760 camera & readying the camera for EVA-3, and
Initiating recharge of the EMU (Extravehicular Mobility Unit) batteries in the A/L BSA (Battery Stowage Assembly.
Kononenko switched the Russian Vozdukh CO2 removal system to manual Mode 5 via the on-board computer system. (Mode 5 uses two adsorbent cartridges with 10 min cycle time, a vacuum pump cycle time of 1 min, plus 40% airflow during sleep, 60-80% during the day, and 100% during physical exercise, for 3-4 crewmembers.)

In the SM, Volkov completed the routine maintenance of the SOZh/ECLSS system, including ASU toilet facilities systems/replaceables (see below for update). (Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier.)

The CDR also performed the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

FE-2-17 Chamitoff spent more time (~15 min.) on familiarizing himself with the onboard CMS (Crew Medical Systems) exercise equipment.

In addition, Greg had another hour for himself for general orientation (station familiarization & acclimatization) as is standard daily rule for fresh crewmembers for the first two weeks after starting station residency.

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1, FE-2), RED resistive exercise device (CDR, FE-1) and VELO cycle with bungee cord load trainer (CDR, FE-1).

Afterwards, Sergey copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

At ~8:15pm EDT, Chamitoff is scheduled for another regular PMC (Private Medical Conference) via S- & Ku-band audio/video.

CDRA Update: After yesterday's R&R of its #2 desiccant/sorbent (#202), the Carbon Dioxide Removal Assembly was declared operational. The fan speed was commanded to 130,000 RPM (revolutions per minute) and will remain at that speed for the remainder of the docked mission.

ASU Update: After the replacement of the failed MNR-NS gas/liquid separator pump of the Russian ASU toilet system by FE-1 Kononenko yesterday, the crew performed several flushes with water to verify a successful R&R, after which ground specialists declared the ASU 'Go' for use. Prior to replacement of the pump, the crew had been using the backup Wring Receptacle Assembly instead of the ASU. The crew requested instructions on how to package the failed pump on Flight 1J.

No CEO (Crew Earth Observation) photo targets uplinked for today.

More... - Chronology...

Home - Browse - Contact
© / Conditions for Use